# TabTreeFormer: Tabular Data Generation Using Hybrid Tree-Transformer

### v2

@@ Abstract
Transformers have achieved remarkable success in tabular data generation. However, they lack domain-specific inductive biases which are critical to preserving the intrinsic characteristics of tabular data. Meanwhile, they suffer from poor scalability and efficiency due to quadratic computational complexity. In this paper, we propose TabTreeFormer, a hybrid transformer architecture that incorporates a tree-based model that retains tabular-specific inductive biases of non-smooth and potentially low-correlated patterns due to its discreteness and non-rotational invariance, and hence enhances the fidelity and utility of synthetic data. In addition, we devise a dual-quantization tokenizer to capture the multimodal continuous distribution and further facilitate the learning of numerical value distribution. Moreover, our proposed tokenizer reduces the vocabulary size and sequence length due to the limited dimension-wise semantic meaning and training set size of tabular data, rendering a significant model size shrink without sacrificing the capability of the transformer model. We evaluate TabTreeFormer on 10 datasets against multiple generative models on various metrics; our experimental results show that TabTreeFormer achieves superior fidelity, utility, privacy, and efficiency. Our best model yields a 40% utility improvement with 1/16 of the baseline model size.
@@ 1 Introduction
Figure 1: Performance comparison of TabTreeFormer (Ours) and other representative tabular generative models in terms of fidelity, utility (i.e., MLE), privacy, and efficiency. TabTreeFormer achieves the best overall performance and outperforms other models on almost all sole metrics.
Tables are one of the most prevalent data modalities in real-world applications (healthcare (Dash et al., 2019), financial services (Assefa et al., 2021), etc.), yet are heavily under-exploited due to privacy concerns (EU, 2016). Fortunately, synthetic data offers an alternative to the utilization of tabular data by modeling the characteristics of real data and reducing the risk of data breach (PDPC, 2024), thus drawing considerable attention in recent years.
The state-of-the-art research shows that transformers such as autoregressive transformers (Borisov et al., 2023), masked transformers (Gulati & Roysdon, 2024), and diffusion models with transformers (Zhang et al., 2024) have achieved impressive fidelity, utility, and privacy in tabular data generation, enabling synthetic data to empower a variety of fields (Hernandez et al., 2022; Assefa et al., 2021). However, unlike the application of transformers in other research areas such as computer vision (Dosovitskiy et al., 2021) and natural language processing (Vaswani et al., 2017), existing transformer models for synthetic tabular data often overlook domain-specific priors (i.e., inductive biases). For example, CvT (Wu et al., 2021) introduces convolutional embedding and projection to transformers to boost vision performances, and Transformer-XL (Dai et al., 2019) introduces segment-level recurrence to transformers to capture longer-term dependencies in long text. While vision and language models have enjoyed the performance boost introduced by domain-specific priors, less exploration has been conducted to leverage them in tabular generative models. Moreover, transformers suffer from poor scalability due to quadratic computational complexity. This raises two interesting questions:
@@
    Are tabular-specific inductive biases beneficial to the performance of tabular generative models?
@@
    Can one exploit unique characteristics of tabular data to improve model scalability and efficiency?
To answer the questions, we propose TabTreeFormer, a generative hybrid transformer that incorporates a tree-based model and a deliberately devised tokenizer to process and introduce tabular-specific inductive biases. We also leverage the limited per-dimension semantic meaning and limited training set size of typical tabular data to enable drastically reduced model size by our tokenizer design.
Inspired by the superior performance of tree-based models on tabular tasks (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2024), we propose to employ them on capturing tabular-specific inductive biases of non-smoothness and potentially low-correlation (Grinsztajn et al., 2024). Non-smoothness is attributed to the existence of discrete features and non-smooth relations between discrete and/or continuous features. Neural networks are biased to overly smooth solutions while tree-based models are exceptionally good at capturing the non-smooth patterns since they learn piece-wise constant functions rather than low-frequency functions (i.e., the functions neural networks tend to bias toward) (Rahaman et al., 2019). Low-correlated features includes uninformative features that contribute trivially to downstream tasks such as classification and regression. Due to the non-rotationally invariant nature of tree-based models (Ng, 2004), trees are more robust against low-correlated features, while neural networks are biased towards the universal presence of correlation.
Nevertheless, trees do not capture multimodal distribution in continuous features, which is a tabular inductive bias particularly essential for generative tasks (Xu et al., 2019; Shi et al., 2024). We then propose dual-quantization tokenization of numeric values, where the first quantization is a K-Means (MacQueen, 1967) quantization that captures multimodal distribution, and the second quantization is a quantile quantization that is used to present the precise value.
Considering the structured nature of tabular data and limited semantic meaning in each dimension, we further optimize the tokenizer so that the vocabulary size and sequence length can be reduced significantly. Together with the limited of training data compared to vision or language tasks, this leads to room for reducing other hyperparameters such as hidden dimensions, and eventually the entire model size.
@@
  • To the best of our knowledge, we are the first work that introduces tabular-specific inductive biases to improve the performance of tabular generative models. We propose a novel hybrid TFM coupled with a TBM to capture such inductive biases.
  • We devise a dual-quantization tokenizer that models the multimodal distribution of continuous values in real data. This practice facilitates the learning of the numerical distribution and further improves the quality of synthetic tabular data.
@@
  • We evaluate TabTreeFormer across 10 datasets against multiple representative tabular generative models and show that TabTreeFormer outperforms other models in fidelity, utility, privacy, and efficiency, as shown in Fig. 1. Notably, with just 1/16 of the baseline model size, our best model achieves a 40% utility improvement over the best-performing baseline model.
@@ 2.1 Tabular Data Generation
Early works on tabular data generation predominantly utilized MLPs and CNNs as backbone architectures and GAN and VAE as generation methods (Park et al., 2018; Xu et al., 2019; Zhao et al., 2021, 2024). Recent works evolved to use transformers with auto-regression, masked modeling, and diffusion as generative paradigms (Borisov et al., 2023; Zhang et al., 2024; Gulati & Roysdon, 2024). For example, TabMT (Gulati & Roysdon, 2024) employs a masked transformer with ordered embedding and achieves outstanding utility in downstream tasks with improved scalability, and TabSyn (Zhang et al., 2024) encodes tabular data into a latent space and generates tabular data using diffusion model, demonstrating impressive fidelity and utility. Despite their enhanced performance, these works overlook the essential inductive biases such as non-smoothness and low-correlated features that are crucial to preserving the intrinsic characteristics of tabular data.
Xu et al. (2019) finds that multimodal distribution in continuous features is a critical inductive bias for table generation. CTGAN (Xu et al., 2019) uses variational Gaussian mixture model (Bishop, 2006) to decompose multimodal values while TabDiff (Shi et al., 2024) introduces a multimodal stochastic sampler. In this work, we introduce multimodal distribution, into an auto-regressive transformer, which was absent in prior works.
@@ 2.2 Tree-based Models for Tabular Data
In recent years, tree-based gradient boosting algorithms have emerged as the dominant approach for tabular data tasks, particularly with the introduction of models like XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018). These tree-based models are ensembles of small and intuitive decision trees that naturally understand feature relations and interactions, thus achieving high predictive performance, efficiency, and scalability, making them the preferred choice in both academia and industry. While some recent research has shown promise of transformers in tabular data (Huang et al., 2020; Arik & Pfister, 2021; Hollmann et al., 2023) in certain scenarios, their computational cost and suboptimal performance make them still less favorable compared to tree-based models (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2024), which is attributed to tabular-specific inductive biases, including non-smoothness and potentially low-correlation, captured by tree-based models (Grinsztajn et al., 2024). In this work, we leverage these features of tree-based models and retain tabular-specific inductive bias to facilitate the training of our transformer model.
@@ 2.3 Tokenizers for Tabular Data
Transformers are not originally designed for tabular data, entailing a need for a tokenizer to translate tabular data to transformer-friendly input tokens or embeddings. Arik & Pfister (2021); Zhang et al. (2024) translate tabular data into a continuous space by completely reformulating the tokenization and embedding layers. Borisov et al. (2023); Solatorio & Dupriez (2023); Gulati & Roysdon (2024) translate tabular data to a sequence of tokens (i.e., discrete space) by a tabular-specific tokenization method, and retain the embedding layer with moderate modifications, which is akin to natural language and more suitable for auto-regressive and masked transformers. In this paper, we opt for the second choice, to optimize the tokenizer for tabular data.
3 TabTreeFormer
Figure 2: Overview of TabTreeFormer. Tokens representing tree leaves are prepended to each row’s tokenized representation, which is also heavily optimized. Tree leaves and optimized data tokens introduce essential inductive biases to the tabular data generator.
Recall that our goal is to improve the performance and efficiency of the generative model by introducing domain-specific inductive biases and shrinking the model size. Formally, let $\mathbf{X}$ be the tabular training dataset for a generative model with $n$ rows, $m_{d}$ discrete and $m_{c}$ continuous features, we aim to generate synthetic tabular data $\mathbf{X}^{\prime}$ similar to $\mathbf{X}$. To achieve this, we propose TabTreeFormer that consists of three building blocks: a tree-based model, a tokenizer, and a transformer model, as shown in Fig. 2. The tree-based model captures some tabular-specific inductive biases, and the tokenizer is used to grasp the multimodal distribution while reducing vocabulary size and sequence length simultaneously. The transformer model learns the prior extracted from the tree-based model and tokenizer and produces high-quality synthetic data. Concrete accounts of training and generation algorithms are described in Algorithm 1 and 2.
@@ 3 TabTreeFormer
2:  Output: Fitted tree-based model $\mathcal{T}$ and the leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, and transformer-based language model $\mathcal{G}$
@@
1:  Input: Leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, transformer-based language model $\mathcal{G}$, and number of rows to be generated $n^{\prime}$
@@ 3.1 Tree-based Models
To introduce tabular-specific inductive biases, we incorporate tree-based models into the transformer (“Tree-based Model” in Fig. 2). This is achieved by augmenting each token sequence representing one table row with additional information derived from the leaf indices of a tree-based model. Formally, let the fitted tree-based model be $\mathcal{T}$ with $T$ trees. Let $l_{i}\in\mathbb{N}$ be the number of leaves in $i$-th tree, where $i\in\{1,2,\dots,T\}$. On $\mathcal{T}$, we have the leaf index matrix $\mathbf{J}=[j_{ik}]\in\mathbb{N}^{n\times T}$ of $\mathbf{X}$ where $j_{ik}\in\{1,2,\dots,l_{k}\}$ represents the leaf index of $i$-th row $\mathbf{X}_{i}$ in the $k$-th tree. Leaf indices provide flattened positional information of each row in a tree, and hence inherently introduce tabular inductive biases captured by trees – non-smoothness and potentially low-correlations. With leaf indices prepended to actual data values, although the subsequent step is still a neural network, tabular inductive biases from tree-based models are inherited by the entire synthetic data generation model through the construction of the intermediate input sequence of tokens, thereby improving the quality of the generated synthetic data. Leaf indices are also provided as prompts during generation, allowing indirect yet informative hints to generate more realistic data.
@@ 3.2 Tokenizer
To model the multimodal distribution of continuous values and reduce tokenized value sizes to eventually reduce the model size, we devise a tailored data tokenizer for tabular data (“Tokenizer” in Fig. 2). For each continuous feature, we quantize the values so that it is of a more natural data format for transformers to learn. However, unlike TabMT (Gulati & Roysdon, 2024) that uses a single quantizer, we use a dual-quantizer (i.e., double quantizers) to capture multimodal continuous distribution and describe the precise values respectively. Transformers do not expect near-Gaussian-distributed input tokens, so we capture the multimodal distribution simply by a K-Means quantizer (MacQueen, 1967) with a small number of clusters (e.g. $K=10$). Following that, the original value is then quantized into a large number of quantiles (e.g. $Q=1000$), based on which original values are recovered, and thus the values are highly precise. By the dual-quantization tokenization method, each numeric value are encoded into 2 discrete values: a bin ID in the K-Means quantizer, and a quantile ID in the quantile quantizer, making 2 tokens sufficient to represent a numeric value with useful and sufficient information. Categorical features are simply tokenized by label encoding, requiring 1 token. Formally, a reversible data tokenizer $\mathcal{D}$ can be fitted on $\mathbf{X}$. Let $c_{i}\in\mathbb{N},b_{i}\in\{0,1,\dots,K\},q_{i}\in\{0,1,\dots,Q\}$ denote the number of categories, K-Means bins, and quantiles per column respectively, where $c_{i}=0$ in continuous features and $b_{i}=q_{i}=0$ in categorical features. Then, the domain for a categorical value is $\{1,2,\dots,c_{i}\}\subseteq\mathbb{N}_{+}$, and the domain for a continuous value is $\{1,2,\dots,b_{i}\}\times\{1,2,\dots,q_{i}\}\subseteq\mathbb{N}_{+}^{2}$.

### v5

@@ Abstract
Transformers have achieved remarkable success in tabular data generation. However, they lack domain-specific inductive biases which are critical to preserving the intrinsic characteristics of tabular data. Meanwhile, they suffer from poor scalability and efficiency due to quadratic computational complexity. In this paper, we propose TabTreeFormer, a hybrid transformer architecture that incorporates inductive biases of tree-based models (i.e., non-smoothness and non-rotational invariance) to process the potentially discrete and low-correlated patterns in tabular data, and hence enhances the fidelity and utility of synthetic data. In addition, we devise a novel tokenizer to capture the multimodal continuous distribution and further facilitate the learning of numerical value distribution. Moreover, our proposed tokenizer reduces the vocabulary size and sequence length based on the complexity of tabular data, shrinking the model size significantly without sacrificing the capability of the transformer model. We evaluate TabTreeFormer on 10 datasets against multiple generative models in various metrics; our experimental results show that TabTreeFormer achieves superior fidelity, utility, privacy, and efficiency. Our best model yields a 40% utility improvement with 1/16 of the baseline model size. Code is available at: https://anonymous.4open.science/r/tabtreeformer-9585.
@@ 1 Introduction
Figure 1: Performance comparison of TabTreeFormer (Ours) and other representative tabular generative models in terms of fidelity, utility (i.e., MLE), privacy, and efficiency. TabTreeFormer outperforms other models on almost all metrics.
Tables are one of the most prevalent data modalities in real-world applications (e.g., healthcare (Dash et al., 2019), financial services (Assefa et al., 2021), etc.), yet are heavily under-exploited due to privacy concerns (EU, 2016). Fortunately, synthetic data offers an alternative to the utilization of tabular data by modeling the characteristics of real data and reducing the risk of data breach (PDPC, 2024), thus drawing considerable attention in recent years.
State-of-the-art research shows that transformers such as autoregressive transformers (Borisov et al., 2023), masked transformers (Gulati & Roysdon, 2023), and diffusion models with transformers (Zhang et al., 2024) have achieved impressive fidelity, utility, and privacy in tabular data generation, allowing synthetic data to empower a variety of fields (Hernandez et al., 2022; Assefa et al., 2021). However, unlike the application of transformers in other research areas such as computer vision (Dosovitskiy et al., 2021) and natural language processing (Vaswani et al., 2017), existing transformer models for synthetic tabular data often overlook domain-specific priors (i.e., inductive biases). For example, CvT (Wu et al., 2021) introduces convolutional embedding and projection to transformers to boost vision performances, and Transformer-XL (Dai et al., 2019) introduces segment-level recurrence to transformers to capture longer-term dependencies in long text. While vision and language models have enjoyed the performance boost introduced by domain-specific priors, less exploration has been conducted to leverage them in tabular generative models. Moreover, transformers suffer from poor scalability due to quadratic computational complexity. This raises two interesting questions:
@@
    What inductive biases are beneficial to the fidelity and utility of synthetically generated tabular data?
@@
    How can one exploit unique characteristics of tabular data to improve model scalability and efficiency?
To answer the questions, we propose TabTreeFormer, a generative hybrid transformer incorporating a tree-based model and a deliberately devised tokenizer to process and introduce tabular-specific inductive biases. We also leverage the limited per-dimension semantic meaning (i.e., each dimension corresponds to at most one feature) of typical tabular data to enable a significant reduction in model size by our tokenizer design.
Tree-based models excel at tabular classification and regression tasks (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2021), which is attributed to their inductive biases to capture tabular characteristics such as non-smoothness and potentially low-correlation (Grinsztajn et al., 2022). Non-smoothness is due to the existence of discrete features and non-smooth relations between discrete and/or continuous features. Tree-based models learn piece-wise constant functions rather than low-frequency functions (i.e., the functions neural networks tend to learn) (Rahaman et al., 2019), so that they are exceptionally good at capturing such non-smooth patterns. textitLow-correlated ifeatures nclude uninformative features that contribute trivially to downstream tasks such as classification and regression. Due to the non-rotationally invariant nature of tree-based models (Ng, 2004), trees are more robust against low-correlated features, while neural networks are biased towards stronger correlations. Inspired by the success of tree models on tabular tasks due to their inductive bias, we propose to employ such priors to facilitate the performance of tabular generative tasks.
However, tree models are not panaceas as they do not inherently capture multimodal distribution in continuous features (i.e., features with probability density functions possessing multiple modes/peaks), which is an essential challenge for tabular generative tasks. Inspired by the prior work (Xu et al., 2019) that tackles the challenge using multimodal decomposition, we propose dual-quantization tokenization for numeric values, a variant of multimodal decomposition catered for transformers where the first quantization is a K-Means (MacQueen, 1967) quantization that captures the multimodal distribution, and the second quantization is a quantile quantization for more precise representation of numerical values.
Moreover, we further optimize the tokenizer to significantly reduce vocabulary size and sequence length by considering the structured nature of tabular data and limited semantic meaning in each dimension (i.e., sole feature for each dimension). In addition, the limited complexity of training data in tables (compared to vision and language tasks) leads to room for reducing model complexity (e.g., hidden dimensions, number of heads, number of layers, etc.) and eventually slimming down the entire model size.
@@
  • To the best of our knowledge, we are the first work that introduces tabular-specific inductive biases to transformers to improve tabular generative models. We propose a novel hybrid transformer coupled with a tree-based model to bring in such inductive biases.
  • We devise a dual-quantization tokenizer that models the multimodal distribution of continuous values. This practice facilitates learning of numerical distribution and further improves the quality of synthetic data.
@@
  • We evaluate TabTreeFormer across 10 datasets against 8 representative tabular generative models and show that TabTreeFormer outperforms other models in fidelity, utility, privacy, and efficiency, as shown in Fig. 1. Notably, with just 1/16 of the baseline model size, our best model achieves a 40% utility improvement over the best-performing baseline model.
@@ 2.1 Tabular Data Generation
Early works on tabular data generation predominantly utilized MLPs and CNNs as backbone architectures and GAN and VAE as generation methods (Park et al., 2018; Xu et al., 2019; Zhao et al., 2021, 2024). Recent works evolved to use transformers with auto-regression, masked modeling, and diffusion as generative paradigms (Borisov et al., 2023; Zhang et al., 2024; Gulati & Roysdon, 2023). For example, TabMT (Gulati & Roysdon, 2023) employs a masked transformer with ordered embedding and achieves outstanding utility in downstream tasks with improved scalability, and TabSyn (Zhang et al., 2024) encodes tabular data into a latent space and generates tabular data using diffusion model, demonstrating impressive fidelity and utility. Despite their enhanced performance, these works overlook the essential inductive biases to capture non-smoothness and low-correlated features that are crucial to preserving the intrinsic characteristics of tabular data.
Xu et al. (2019) finds that multimodal distribution in continuous features is a critical aspect for tabular generative models’ inductive bias to capture. CTGAN (Xu et al., 2019) uses variational Gaussian mixture model (Bishop, 2006) to decompose multimodal values while TabDiff (Shi et al., 2024) introduces a multimodal stochastic sampler. In this work, we devise a tokenizer that models the multimodal distribution and introduces the prior into the auto-regressive transformer.
@@ 2.2 Tree-based Models for Tabular Data
Tree-based gradient boosting algorithms are the dominant approach for tabular tasks, notable examples are XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018). These tree-based models are ensembles of small and intuitive decision trees that naturally understand feature relations and interactions, thus achieving high predictive performance, efficiency, and scalability. Although recent research has shown promise of transformers in tabular data-related tasks (Huang et al., 2020; Arik & Pfister, 2021; Hollmann et al., 2023), their suboptimal performance caused by the absence of tabular-specific inductive biases makes them still less favorable to tree-based models in many scenarios (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2021). Meanwhile, tree-based generative models (Watson et al., 2023; McCarter, 2024), while being efficient, still fall short of synthetic data quality compared to deep neural networks like transformers. In this work, we take advantage of both models and employ tree-based models to introduce tabular-specific inductive bias and facilitate the learning of the transformer model.
@@ 2.3 Tokenizers for Tabular Data
Transformers are not originally designed for tabular data, entailing a need for a tokenizer to translate tabular data to transformer-friendly input tokens or embeddings. Arik & Pfister (2021); Zhang et al. (2024) translate tabular data into a continuous space by completely reformulating the tokenization and embedding layers. Borisov et al. (2023); Solatorio & Dupriez (2023); Gulati & Roysdon (2023) translate tabular data to a sequence of tokens (i.e., discrete space) by a tabular-specific tokenization method, and retain the embedding layer with moderate modifications, which is akin to natural language and more suitable for transformers. In this paper, we opt for the second choice to optimize the tokenizer for tabular data.
@@
2:  Output: Fitted tree-based model $\mathcal{T}$ and the leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, and transformer $\mathcal{G}$
@@
3 TabTreeFormer
Figure 2: Overview of TabTreeFormer (data flow from bottom to top). TabTreeFormer consists of 3 components: i) a tree-based model that introduces tabular-specific inductive biases; ii) a tokenizer that efficiently and compactly represents data while capturing multimodal distributions; iii) a transformer model that learns the priors extracted from the tree and tokenizer to generate high-quality synthetic data.
Recall that our goal is to improve the performance and efficiency of the generative model by introducing domain-specific inductive biases and shrinking the model size. Formally, let $\mathbf{X}$ be the tabular training dataset for a generative model with $n$ rows, $m_{d}$ discrete and $m_{c}$ continuous features, we aim to generate synthetic tabular data $\mathbf{X}^{\prime}$ similar to $\mathbf{X}$. To achieve this, we propose TabTreeFormer that consists of three building blocks: a tree-based model, a tokenizer, and a transformer model, as shown in Fig. 2. The tree-based model captures some tabular-specific inductive biases, and the tokenizer is used to grasp the multimodal distribution while reducing vocabulary size and sequence length simultaneously. The transformer model learns the priors extracted from the tree-based model and tokenizer to produce high-quality synthetic data. Concrete accounts of training and generation algorithms are described in Algorithm 1 and 2.
@@ 3 TabTreeFormer
1:  Input: Leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, transformer-based language model $\mathcal{G}$, and number of rows generate $n^{\prime}$
@@ 3.1 Tree-based Models
To introduce tabular-specific inductive biases, we incorporate tree-based models into the transformer (“Tree-based Model” in Fig. 2). This is achieved by augmenting each token sequence representing one table row with additional information derived from the leaf indices of a tree-based model fitted on the prediction task defined on the dataset or a random feature if such a task is unavailable. Formally, let the fitted tree-based model be $\mathcal{T}$ with $T$ trees. Let $l_{i}\in\mathbb{N}$ be the number of leaves in $i$-th tree, where $i\in\{1,2,\dots,T\}$. In tree $\mathcal{T}$, there is a leaf index matrix $\mathbf{J}=[j_{ik}]\in\mathbb{N}^{n\times T}$ of $\mathbf{X}$ where $j_{ik}\in\{1,2,\dots,l_{k}\}$ represents the leaf index of $i$-th row $\mathbf{X}_{i}$ in the $k$-th tree. Leaf indices provide flattened positional information of each row in a tree, and hence inherently introduce the non-smooth and non-rotationally invariant inductive biases of trees to capture the non-smoothness and potentially low-correlations of tabular data. With leaf indices prepended to actual data values, although the subsequent model is still a neural network, tabular-specific inductive biases from tree-based models are inherited by the entire generative model through the construction of the intermediate input token sequence (i.e., indicated by the red arrows in Fig. 2), thereby improving the quality of the synthetic data. Leaf indices are also served as prompts during inference, allowing indirect yet informative hints to generate more realistic data.
@@ 3.2 Tokenizer
To model the multimodal distribution of continuous values and reduce the vocabulary size along with the tokenized input sequence length to eventually reduce the model size, we devise a tailored data tokenizer for tabular data (“Tokenizer” in Fig. 2). For each continuous feature, we quantize the values so that it is of a more natural data format for transformers to learn. However, unlike TabMT (Gulati & Roysdon, 2023) that uses a single quantizer, we use a dual-quantizer (i.e., double quantizers) to capture multimodal continuous distribution and describe the precise values respectively. Transformers do not expect near-Gaussian-distributed input tokens, so we capture the multimodal distribution simply by a K-Means quantizer (MacQueen, 1967) with a small number of clusters (e.g., $K=10$). Following that, the original value is then quantized into a large number of quantiles (e.g., $Q=1000$), based on which original values are recovered, and thus the values are highly precise.
Employing the dual-quantization tokenization scheme, each numeric value is encoded into two discrete values: a bin ID in the K-Means quantizer and a quantile ID in the quantile quantizer, allowing two tokens to represent a numeric value with valuable and sufficient information. Categorical features are simply tokenized by label encoding, requiring 1 token. Formally, a reversible data tokenizer $\mathcal{D}$ can be fitted on $\mathbf{X}$. Let $c_{i}\in\mathbb{N},b_{i}\in\{0,1,\dots,K\},q_{i}\in\{0,1,\dots,Q\}$ denote the number of categories, K-Means bins, and quantiles per column respectively, where $c_{i}=0$ in continuous features and $b_{i}=q_{i}=0$ in categorical features. Then, the domain for a categorical value is $\{1,2,\dots,c_{i}\}\subseteq\mathbb{N}^{+}$, and the domain for a continuous value is $\{1,2,\dots,b_{i}\}\times\{1,2,\dots,q_{i}\}\subseteq(\mathbb{N}^{+})^{2}$.